Data was collected for 370 randomly selected 10 minute intervals. For each ten-minute interval, the number of people entering the atrium of a large mall were recorded. The data is summarized in the table below.

Number
  of Guests  
   Frequency   
40 – 49 33
50 – 59 93
60 – 69 76
70 – 79 76
80 – 89 92



What is the class width for this Frequency Distribution Table?

Class width =
Answer #1

Ans ) 10

Explanation: There are two ways to find the class width.

First make number of guests in continuous data by subtracting 0.5 from lower limit and adding 0.5 in upper limit.

39.5 - 49.5

49.5 - 59.5

59.5 - 69.5
69.5 - 79.5
79.5 -89.5

Class width = upper limit - lower limit

= 49.5- 39.5 = 10 ans

Another way subtract lower limit of two consecutive row either lower or upper limit.

For example

50-40 = 10 ans